Iron has a specific gravity of 7.8. Suppose you have a solid cube-shaped piece of iron and know that its mass is 56 kg. The volume of a cube with edge a is
1
Expert's answer
2019-05-17T11:50:08-0400

The volume of iron cube

"V=\frac{m}{\rho}=\frac{m}{1000\:\rm{SG}}"

"=\frac{56\:\rm{kg}}{1000\times 7.8\:\rm{kg/m^3}}=7.179\times 10^{-3}\:\rm{m^3}"